OpenStudy (anonymous):

Which expression is equivalent to (1 + 9x)10 ? A. –10 – 90x B. 11 + 9x C. 10 + 90x D. –11 – 9x

jigglypuff314 (jigglypuff314):

distributive property: a(b+c) = ab + ac 10(1+9x) = 10*1 + 10*9x = 10 + 90x